OpenStudy (amonoconnor):

I have a brief conceptual question! When evaluating Volumes of Solid of Revolution, and the following mathematical circumstance arises, within the integral... 4 - 4x^(-1) ... Which of the following should it be rewritten as? 1. 4 - 4(ln(x)) 2. 4 - ln(4x) 3. Or something completely different(?) Any and all help is greatly appreciated! :)

OpenStudy (anonymous):

\[\int\limits_{}^{}4-4x^{-1}\]\[\int\limits_{}^{}4-\frac{ 4 }{ x }\]\[\int\limits_{}^{}4-4\int\limits_{}^{}\frac{ 1 }{ x }\]\[4x-4\int\limits_{}^{}\frac{ 1 }{ x }\]\[4x-4(\ln (x))+C\] or:\[4(x-\ln (x))+C\] ANSWER:\[4x-4(\ln(x) )+C\]

OpenStudy (amonoconnor):

So, if there is a coefficient to the variable that is to the negative first– therefore meaning the anti-derivative of said variable is "undefined", or 0– you detatch the coefficient from the variable when rewriting the term(s) as a natural log, taking the log of only the variable itself (with the given number or limit plugged in of course)??

OpenStudy (amonoconnor):

Simply put, and when analyzing THIS situation only, is "4 - 4x^(-1)" equal to "4 - 4(ln(x))" or "4 - ln(4x)" ? Thank you!

OpenStudy (amonoconnor):

@ganeshie8 Help Please?!? :D

ganeshie8 (ganeshie8):

\[\int 4 - \dfrac{4}{x} dx = 4x - 4\ln |x| + C \]

ganeshie8 (ganeshie8):

we're done with integration/calculus part.

ganeshie8 (ganeshie8):

if you prefer, u can write the answer as : \[ 4x - 4\ln |x| + C = 4x - \ln |x|^4 + C \]

ganeshie8 (ganeshie8):

^^thats just based on a log property : \[m \log n = \log n^m\]

ganeshie8 (ganeshie8):

I'm not sure if i have answered ur question...

OpenStudy (amonoconnor):

Yes, you have! Thank you very much!

ganeshie8 (ganeshie8):

np :)

ganeshie8 (ganeshie8):

https://www.youtube.com/watch?v=iNtMLGvzFHA

ganeshie8 (ganeshie8):

i think u may like watching that funny debate on derivatives Vs integrals... when u have time :)

OpenStudy (amonoconnor):

Alright, I'll check it out;)